OpenStudy (anonymous):

Write a polynomial that would have 3 turning points, shifted to the right 7 units and be reflected over the x-axis.

OpenStudy (triciaal):

a = -1 reflection shift to the right x -7 3 turning points highest power in polynomial 3 f(x) = -(x-7)^3

OpenStudy (triciaal):

this graph only shows 2 turns when plugged into the graphing calculator. therefore not sure

OpenStudy (triciaal):

@DanJS is this incomplete? @sleepyjess ?

OpenStudy (danjs):

a polynomial of 4th degree will have 3 or less turning points

OpenStudy (danjs):

where there is a local max or min

OpenStudy (triciaal):

did you see what this graph look like?

OpenStudy (danjs):

ie - x^4 - 5x^2

OpenStudy (triciaal):

would it be Ok to just add a factor x?

OpenStudy (triciaal):

normally take the 2nd derivative to find the max or min

OpenStudy (triciaal):

so what would be appropriate for this? something we make up or is it specific with the information given?

OpenStudy (danjs):

hmm, try factors i guess. y = -(x-7)(x-3)(x+3)

OpenStudy (danjs):

that is a 3rd degree derivative

OpenStudy (triciaal):

this worked for me according to the graph -x^2(x-7)^2

OpenStudy (danjs):

y = (x+3)(x-3)(x+7)(x-7) and let x=x-7 for the shift y = -((x-7)+7)((x-7)-7)*((x-7)+3)*((x-7)-3)

OpenStudy (triciaal):

the one above max(0,0) min (3.5, -150) max (7,0)

OpenStudy (triciaal):

where did you get +/-3 to use as factor?

OpenStudy (danjs):

try y = -(x-8)^2 * (x-6)^2

OpenStudy (danjs):

or

OpenStudy (triciaal):

https://www.desmos.com/calculator

OpenStudy (danjs):

y = -(x-5)^2*(x-9)^2

OpenStudy (danjs):

y = -(x - 4)^2*(x-10)^2

OpenStudy (danjs):

notice the pattern

OpenStudy (triciaal):

@jazzyrocroyal any ideas now?

OpenStudy (danjs):

-(x-3)^2(x-11)^2

OpenStudy (danjs):

so on...

OpenStudy (danjs):

this is just from me playing around. lol there may be a formal way to do that that is slipping my mind

OpenStudy (triciaal):

@DanJS thanks for your help. @jazzyrocroyal still nothing to say? @Directrix hi just in time to close this out for us

OpenStudy (danjs):

welcome, i would just go with y = -(x-8)^2 * (x-6)^2

OpenStudy (danjs):

If you have a -x^4 and a +x^2 term it works -x^4 + x^2 -(x-7)^4 + (x-7)^2

OpenStudy (danjs):

that seems to be the trend with them, in general a 4th order polynomial will have 3 OR LESS turns, to have 3, looks like you need Opposite signs on the 4th order and 2nd order terms

OpenStudy (danjs):

and to have one opening down< you want a negative on the 4th order term, so you need a positive 2nd order term

OpenStudy (triciaal):

exploring some more 8^2 +6^2 = 100 (8^2*6^2 )1/2 = 48 100 + 48 = 148 for y then only use the negative close enough to -150 (8*6)^1/2 then divide2 = 3.4 no use 7/2 = 3.5 the x coordinate for the vertex

OpenStudy (triciaal):

just a sample? this is same as @DanJS basically in regards to the choosing of exponents

OpenStudy (triciaal):

will it satisfy the requirement for this or just a 3-point graph?

OpenStudy (triciaal):

I looked. that's just a sample.

OpenStudy (triciaal):

repost this worked for me according to the graph -x^2(x-7)^2

OpenStudy (triciaal):

did you look at the graph?

OpenStudy (danjs):

Think you just need to alternate the signs on each decreasing power of x, with the first being -x^4 so it points down, then replacing x by(x-7) for the shift

OpenStudy (triciaal):

@jazzyrocroyal feel free to join in especially since this is your post

OpenStudy (danjs):

all even powers will be negative, all odd powers positive coefficients

OpenStudy (danjs):

does this have something to do with Pascal's triangle, i vaguely remember doing that

OpenStudy (danjs):

ah, got to run for the time, good luck ya'll

OpenStudy (triciaal):

@DanJS execellent job thanks